# Improving probabilistic forecasts of extreme wind speeds by training statistical post-processing models with weighted scoring rules

### v2

@@ 8 Discussion and conclusions
In this work we have introduced the usage of threshold-weighted scoring rules to train probabilistic post-processing models, tailored to extremes. This is an alternative to schemes based on weighting loss functions by outcomes which different researchers have been proposing in the literature, which however are problematic as they might encourage "hedging" (Lerch et al., 2017). We have shown the positive effects of training with the twCRPS on forecasts’ extreme event performance. We have identified a distribution body-tail trade-off which manifests as a trade-off between CRPS and twCRPS performance. We have shown that training using a twCRPS is possible for a variety of thresholds of interest and have identified challenges when wanting to optimize sample versions of the twCRPS. We have developed two options for users to balance distribution body and extreme event performance of their forecasts: weighted sums of CRPS and twCRPS and the linear pool between the CRPS and twCRPS forecast. We developed a number of synthetic experiments to show the effect of the twCRPS and finally we have calculated closed-form twCRPS expressions for a number of distributions, which can be found in appendix B together with different characterizations of the twCRPS (appendix A)
@@
There are a number of further research questions that follow from this work and should be explored. First it would be interesting to explore the usage of other weighted scores such as the proper quantile weighted score (Gneiting and Ranjan, 2011) or the proper outcome weighted scores (Holzmann and Klar, 2017) and the usage of different, non-threshold-based, weighting functions when training models with weighted scores with emphasis on different parts of the distribution. Allen et al. (2023b, a) develop multivariate versions of the weighted scores, by showing that weighted scores fall into the class of kernel scores. In order to improve predictions of compound extremes – extremes that can have the multiplicity of the impact of singular extremes (Zscheischler et al., 2020) – investigating the usage of these scores to train models with potentially improved forecasting skill for compound extremes might be of interest. This is especially relevant as for example Chen et al. (2022) show the positive effects of a kernelized loss function for improved deterministic predictions of wind speed extremes and similar approaches could be successful for probabilistic predictions. With more complex loss functions or scores, closed-form solutions of these might not be available, thus it might be of interest to investigate how to stabilize sample based optimization of weighted scores. This might also be relevant when it comes to training or finetuning of generative deep learning models, which have shown great potential for post-processing applications (Chen et al., 2024). As neural networks are able to adjust well to large numbers of covariates, this could be combined with research into what predictors for post-processing models are most relevant to improve extreme event forecasts. Research on this and the usage of (weighted) scoring rules for training generative neural networks is in progress by the authors. In addition it might be of interest to investigate forecast combination methods further, for combining distribution body and distribution tail forecasts. For example one could explore methods based on online-learning (van der Meer et al., 2024) or pointwise combinations (Berrisch and Ziel, 2023) and see whether they can help to alleviate the trade-off between CRPS and twCRPS performance further.
Secondly, much more sophisticated post-processing methods exist. Many of these are extensions of EMOS or still based on a closed-form distributional structure, for example so-called distributional regression networks (DRN, Rasp and Lerch 2018) parametrize the parameters of a predictive distribution as neural networks. The authors are hopeful that the results in this article might also generalize to other EMOS-like model structures, however we leave this for future research.
@@
Acknowledgements: Jakob Wessel was supported during this work by the Engineering and Physical Sciences Research Council (EPSRC) under grant award number 2696930 and acknowledges support by The Alan Turing Institute’s Enrichment Scheme. The authors would like to thank Sam Allen for pointing out the connection between the threshold-weighted CRPS and the CRPS of a censored distribution, as well as input regarding closed-form expressions of the twCRPS. They are grateful to Jamie Kettleborough and Fiona Spuler for helpful discussions of the results and comments on the manuscript draft.
Code and Data availability: The code for this study can be found at https://github.com/jakobwes/Improving-probabilistic-forecasts-of-extreme-wind-speeds. Unfortunately the authors are unable to share the data, however this can be requested from the UK Met Office.
@@ Appendix A Characterizations and properties of the twCRPS
The proof follows from the definitions of ${\tilde{F}}_{\tau}$ and $v_{\tau}(y)$:
@@
The proof follows by considering integration limits in the CRPS and twCRPS definitions and the definition of $v_{\tau}(y)$:
@@
$=\text{twCRPS}_{\tau}(F,y).$ (27)
The twCRPS is stable under censoring at a point below the threshold $\tau$:
@@
The lower-tail twCRPS can be expressed in terms of the CRPS and the upper-tail twCRPS:
@@ B.1.1 Normal distribution
$\text{twCRPS}_{\tau}\left(\Phi,y\right)=-\tau\Phi^{2}(\tau)+v_{\tau}(y)\left[2\Phi(v_{\tau}(y))-1\right]+2\left[\varphi(v_{\tau}(y))-\varphi(\tau)\Phi(\tau)\right]-\frac{1}{\sqrt{\pi}}\left[1-\Phi(\sqrt{2}\tau)\right],$ (32)
Source: This formula is given by Wessel et al. [2024].
B.1.2 Laplace distribution
twCRPS expression:
$\text{twCRPS}_{\tau}\left(F,y\right)=\begin{cases}|v_{\tau}(y)|-\frac{3}{4}+\exp(-|v_{\tau}(y)|)-\frac{1}{8}\exp(2\tau),&\tau\leq 0,\\ v_{\tau}(y)-\tau+\exp(-v_{\tau}(y))+\frac{1}{8}\exp(-2\tau)-\exp(-\tau),&\tau\geq 0,\\ \end{cases}$ (34)
@@ B.2.3 Log-logistic distribution
twCRPS expression for $\tau\geq 0$ and $10<\sigma<$ 1:
@@ B.3.1 Continuous uniform distribution on $[a,b]$
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=(b-a)\,\text{twCRPS}_{(\tau-a)/(b-a)}\left(F,\frac{y-a}{b-a}\right),$ (45)
@@ B.3.2 Generalized Pareto distribution
$\text{twCRPS}_{\tau}\left(F_{\xi,\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{\xi},\frac{y-\mu}{\sigma}\right),$ (46)
@@ B.3.3 Truncated logistic distribution on $[a,b]$
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=&\frac{1}{[F(b)-F(a)]^{2}}\left\{F(\tau)-F(b)+F^{2}(a)\log\frac{F(v_{\tau}(y))}{F(\tau)}\right.\\ &\left.+[1-F(a)]^{2}\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}+F^{2}(b)\log\frac{F(b)}{F(v_{\tau}(y))}+[1-F(b)]^{2}\log\frac{1-F(v_{\tau}(y))}{1-F(b)}\right\},\end{aligned}$
@@
where $F$ is the standard logistic CDF, $F_{a}^{b}$ is the standard logistic CDF truncated to $[a,b]$ with $-\infty<a<b<\infty$, and $F_{a,\mu,\sigma}^{b}(x)=F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
@@ B.4 Some notes on the integral function $I(\alpha,\beta)$
In the derivation of the twCRPS for the log-normal distribution the integral function $I(\alpha,\beta)$ appears (eq.(43)). It has the statistical interpretation as the probability $P(X>\alpha,Y<X+\beta)$ where $X$ and $Y$ are two independent standard normal random variables. The integral function $I$ can be expressed in terms of Owen’s $T$ function [Owen, 1956, 1980],
@@
Alternatively, we propose a fast and stable method for evaluating $I(\alpha,\beta)$ based on a recursive power series expansion, making use of (probabilist’s) Hermite polynomials ${\{H_{n}\}}_{n=0}^{\infty}$. The integral function $I$ has the Taylor series
@@
$C_{n}=$ $-\frac{\alpha\beta}{n}C_{n-1}-\frac{n-3}{(n-1)n}\beta^{2}C_{n-2},\qquad n\geq 3,$ (60)
$I_{n}=$ $C_{n}-\frac{n-2}{2(n-1)n}\beta^{2}I_{n-2},\qquad n\geq 3.$ (61)
@@
Appendix C Reparametrization trick
In section 5.3 we used Adam, a stochastic gradient descent method, to optimize the parameters $\theta$ of the EMOS post-processing model. This requires gradients of the twCRPS-expectation with respect to $\theta$:
$\nabla_{\theta}\ \text{twCRPS}_{\tau}(F_{\theta},y)=\nabla_{\theta}\left(\mathbb{E}_{X\sim F_{\theta}}|v_{\tau}(X)-v_{\tau}(y)|-\frac{1}{2}\mathbb{E}_{X,X^{\prime}\sim F_{\theta}}|v_{\tau}(X)-v_{\tau}(X^{\prime})|\right).$ (64)
This is difficult to compute through a Monte-Carlo estimate as the expectations, of which we compute the gradient, also depend on $\theta$. To address this we use the reparametrization trick [Kingma and Welling, 2022] i.e. reparametrize $X\sim F_{\theta}$ as $X=g(\epsilon,\theta)$ with $\epsilon\sim Z$, so in terms of a base-distribution $Z$ which does not depend on the parameters. This allows to write the gradient of the expectation as expectation of a gradient:
$\nabla_{\theta}\left(\mathbb{E}_{X\sim F_{\theta}}|v_{\tau}(X)-v_{\tau}(y)|-\frac{1}{2}\mathbb{E}_{X,X^{\prime}\sim F_{\theta}}|v_{\tau}(X)-v_{\tau}(X^{\prime})|\right)$ (65)
$=\mathbb{E}_{\epsilon\sim Z}\ \nabla_{\theta}\ |v_{\tau}(g(\epsilon,\theta))-v_{\tau}(y)|-\frac{1}{2}\mathbb{E}_{\epsilon,\epsilon^{\prime}\sim Z}\ \nabla_{\theta}\ |v_{\tau}(g(\epsilon,\theta))-v_{\tau}(g(\epsilon^{\prime},\theta))|,$ (66)
Such a reparametrization is easily possible for the truncated normal and truncated logistic distributions. For example for the truncated normal we sample:
$g(\epsilon,\mu,\sigma)=\Phi_{\mu,\sigma}^{-1}\left(\Phi_{\mu,\sigma}(0)+\left(1-\Phi_{\mu,\sigma}(0)\right)\cdot\epsilon\right)\text{ with }\epsilon\sim U(0,1),$ (67)
here $\Phi_{\mu,\sigma}$ is the Gaussian CDF. We use the automatic differentiation integrated in PyTorch [Paszke et al., 2019] to compute exact gradients.
Appendix D Additional experiments for sample-based optimization
We analysed sample based optimization for the truncated normal and truncated logistic distribution for the 80th percentile threshold. As seen above the choice of optimization algorithm and starting values can have a strong influence. Furthermore the choice of the $x_{0}$ parameter in the empirical twCRPS formulation can have an influence on the optimization (see eq. 13, 14). For $x_{0}\neq\tau$ the chaining function is discontinuous, which might impact the optimization, and Allen et al. [2023b] find that in a multivariate example setting $x_{0}=0$ only makes the twCRPS marginally more discriminative than the CRPS for upper tail mis-specifications. We find that using BFGS or Nelder-Mead optimizers the optimization often fails for $x_{0}=0$, but at least for Nelder-Mead sometimes succeeds for $x_{0}=\tau$. The value of $x_{0}$ seems to have less of an impact on the Adam optimizer, however on a subset of stations the resulting models seemed to be less efficient. The difficulty of the choice of $x_{0}$ also has possible impacts for (optimizing) generalizations of the twCRPS to higher dimensions [Allen et al., 2023a, b] where no canonical point (such as $x_{0}=\tau$ here) exists to map all excluded values to.
Figure 11 shows the twCRPS and CRPS skill score (relative to the CRPS-trained model, obtained through closed-form CRPS optimization) for models obtained through minimization of the closed form twCRPS and sample based optimization for Nelder-Mead and Adam optimizers. Training using sample versions is much less efficient i.e. leads to models with far smaller twCRPS improvement than training using the closed-form expressions. Optimization using Nelder-Mead seems to converge, however the resulting models do not improve in terms of twCRPS (and CRPS) over the CRPS baseline, somewhat independently from the number of samples. When using Adam, a stochastic gradient descent algorithm, and exact gradients, sample based optimization seems to work in some cases. However for both optimization algorithms there is a high sensitivity on the starting values. For Adam, when using the parameters from the closed-form CRPS optimization as starting values (in whose neighborhood the twCRPS parameters lie) then the resulting models seems to improve slightly in terms of twCRPS over the CRPS-trained model, without reaching the level of the closed-form-twCRPS-trained model. However using the default starting values, Adam based optimization does not lead to any improvement and is worse in terms of CRPS and twCRPS than the closed-form solution model. For Nelder-Mead CRPS starting values lead to models with the same skill as the CRPS-trained model, whereas the default starting values lead to models that are slightly worse than the CRPS-trained models in terms of both CRPS and twCRPS.
Figure 11: twCRPS and CRPS skill score (over the CRPS-trained model) for models trained using closed form twCRPS solutions as well as the sample twCRPS using different optimizers.
Appendix E Additional figures
(a) 80th percentile
(b) 90th percentile
Figure 12: 80th (a) and 90th (b) percentile of the training distribution used as thresholds. Average 80th percentile threshold: 6.4 m/s (std: 2.1 m/s), maximum 80th percentile threshold: 11.8 m/s. Average 90th percentile threshold: 8.0 m/s std: 2.4 m/s), maximum 90th percentile threshold: 14.4 m/s.
@@ Appendix E Additional figures
0.8 0.7 0.157 0.153
0.8 1.701 1.624
0.85 2.300 2.094
0.9 2.071 2.009
0.95 1.118 1.091
0.9 0.7 -10.872 -10.928
0.8 -7.571 -7.859
0.85 -4.508 -4.578
0.9 0.936 0.744
0.95 3.885 3.838
Figure 13: twCRPS skill improvement for the truncated normal twCRPS-trained model and the 90th percentile threshold.
Figure 14: Brier skill score (over the CRPS-trained model) as a function of the threshold for twCRPS-trained models trained using different thresholds. The bands represent 95% intervals over stations.
Figure 15: twCRPS and CRPS for the truncated normal distribution for training using log likelihood, twCRPS and weighted combinations CRPS + $\gamma\cdot$ twCRPS for selected values of $\gamma=2,10,20$.
(a) Weighted Training
(b) Linear Pool
Figure 16: twCRPS skill over the CRPS-trained model for a range of evaluation quantiles for a) models trained using weighted combinations with selected values of $\gamma=2,5,10,20$ and b) linear pool-models for selected values of $\lambda=0.2,0.5,0.7,0.9$. Truncated normal predictive distribution and training threshold as 80th percentile of the training dataset. Note: the twCRPS for an evaluation quantile of 0.0 corresponds to the unweighted CRPS.

### v4

@@ 8 Discussion and conclusions
In this work we have introduced the usage of threshold-weighted scoring rules to train probabilistic post-processing models, tailored to extremes. This is an alternative to schemes based on weighting loss functions by outcomes which different researchers have been proposing in the literature, which however are problematic as they might encourage hedging (Lerch et al., 2017). We have shown the positive effects of training with the twCRPS on forecasts’ extreme event performance. We have identified a distribution body-tail trade-off which manifests as a trade-off between CRPS and twCRPS performance. We have shown that training using a twCRPS is possible for a variety of thresholds of interest and have shown that successful training depends on the training datasize used. We have developed two options for users to balance distribution body and extreme event performance of their forecasts: weighted sums of CRPS and twCRPS and the linear pool between the CRPS and twCRPS forecast. We developed a number of synthetic experiments to show the effect of the twCRPS and finally we have calculated closed-form twCRPS expressions for a number of distributions, which can be found in Appendix B together with different characterizations of the twCRPS (Appendix A)
@@
There are a number of further research questions that follow from this work and should be explored. First it would be interesting to explore the usage of other weighted scores such as the proper quantile weighted score (Gneiting and Ranjan, 2011) or the proper outcome weighted scores (Holzmann and Klar, 2017), censored likelihood scores (de Punder et al., 2023) and the usage of different, non-threshold-based, weighting functions when training models with weighted scores with emphasis on different parts of the distribution. Allen et al. (2023b, a) develop multivariate versions of the weighted scores, by showing that weighted scores fall into the class of kernel scores. In order to improve predictions of compound extremes – extremes that can have the multiplicity of the impact of singular extremes (Zscheischler et al., 2020) – investigating the usage of these scores to train models with potentially improved forecasting skill for compound extremes might be of interest. This is especially relevant as for example Chen et al. (2022) show the positive effects of a kernelized loss function for improved deterministic predictions of wind speed extremes and similar approaches could be successful for probabilistic predictions. With more complex loss functions or scores, closed-form solutions of these might not be available, thus it might be of interest to investigate sample-based optimization of weighted scores. This might also be relevant when it comes to training or finetuning of generative deep learning models, which have shown great potential for post-processing applications (Chen et al., 2024). As neural networks are able to adjust well to large numbers of covariates, this could be combined with research into what predictors for post-processing models are most relevant to improve extreme event forecasts. Research on this and the usage of (weighted) scoring rules for training generative neural networks is in progress by the authors. In addition it might be of interest to investigate forecast combination methods further, for combining distribution body and distribution tail forecasts. For example one could explore methods based on online-learning (van der Meer et al., 2024) or pointwise combinations (Berrisch and Ziel, 2023) and see whether they can help to alleviate the trade-off between CRPS and twCRPS performance further. Alternatively, optimization methods such as multi-objective optimization could be explored to see whether they manage to balance the CRPS and weighted scoring rule objectives.
Secondly, much more sophisticated post-processing methods exist. Many of these are extensions of EMOS or still based on a closed-form distributional structure, for example so-called distributional regression networks (DRN, Rasp and Lerch 2018) parametrize the parameters of a predictive distribution as neural networks. The authors are hopeful that the results in this article might also generalize to other EMOS-like model structures, including other predictive and possibly tail-adaptive distributions. However, we leave this for future research.
@@
Acknowledgements: Jakob Wessel was supported during this work by the Engineering and Physical Sciences Research Council (EPSRC) under grant award number 2696930 and acknowledges support by The Alan Turing Institute’s Enrichment Scheme. The authors would like to thank Sam Allen for pointing out the connection between the threshold-weighted CRPS and the CRPS of a censored distribution, as well as input regarding closed-form expressions of the twCRPS. They are grateful to Jamie Kettleborough and Fiona Spuler for helpful discussions of the results and comments on the manuscript draft. Also, they would like to thank the three reviewers for comments and feedback that have substantially improved the manuscript.
Code and Data availability: The code for this study can be found at https://github.com/jakobwes/Improving-probabilistic-forecasts-of-extreme-wind-speeds. Unfortunately, the authors are unable to share the data, however this can be requested from the UK Met Office.
@@ Appendix A Characterizations and properties of the twCRPS
The proof follows from the definitions of ${\tilde{F}}_{\tau}$ and $v_{\tau}$:
@@
The proof follows by considering integration limits in the CRPS and twCRPS definitions and the definition of $v_{\tau}$:
@@
[…]
$=\text{twCRPS}_{\tau}(F,y).$
Property 1: The twCRPS is stable under censoring at a point below the threshold $\tau$.
@@
Property 2: The lower-tail twCRPS can be expressed in terms of the CRPS and the upper-tail twCRPS.
@@ B.1.1 Laplace distribution
$\text{twCRPS}_{\tau}\left(F,y\right)=\begin{cases}|v_{\tau}(y)|-\frac{3}{4}+\exp(-|v_{\tau}(y)|)-\frac{1}{8}\exp(2\tau),&\tau<0,\\ v_{\tau}(y)-\tau+\exp(-v_{\tau}(y))+\frac{1}{8}\exp(-2\tau)-\exp(-\tau),&\tau\geq 0,\\ \end{cases}$ (28)
@@ B.1.2 Logistic distribution
B.1.3 Normal distribution
twCRPS expression:
$\text{twCRPS}_{\tau}\left(\Phi,y\right)=&-\tau\Phi^{2}(\tau)+v_{\tau}(y)\left[2\Phi(v_{\tau}(y))-1\right]\\ &+2\left[\varphi(v_{\tau}(y))-\varphi(\tau)\Phi(\tau)\right]-\frac{1}{\sqrt{\pi}}\left[1-\Phi(\sqrt{2}\tau)\right],\end{aligned}$
Source: This formula is given by Wessel et al. [2024].
@@ B.2.3 Log-logistic distribution
twCRPS expression for $10<\sigma<$ 1 and $\tau\geq 0$:
@@ B.3.1 Continuous uniform distribution on $[a,b]$
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=(b-a)\,\text{twCRPS}_{(\tau-a)/(b-a)}\left(F,\frac{y-a}{b-a}\right),\qquad a\leq\tau\leq b,\qquad a\leq y\leq b,$ (40)
@@ B.3.2 Generalized Pareto distribution
$\text{twCRPS}_{\tau}\left(F_{\xi,\mu,\sigma},y\right)=\sigma\,\text{twCRPS}_{(\tau-\mu)/\sigma}\left(F_{\xi},\frac{y-\mu}{\sigma}\right),\qquad\qquad\tau\geq\mu,$ (41)
@@ B.3.3 Truncated logistic distribution on $[a,b]$
$\text{twCRPS}_{\tau}\left(F_{a}^{b},y\right)=&\frac{1}{[F(b)-F(a)]^{2}}\left\{F(\tau)-F(b)+F^{2}(a)\log\frac{F(v_{\tau}(y))}{F(\tau)}\right.\\ &\left.+[1-F(a)]^{2}\log\frac{1-F(\tau)}{1-F(v_{\tau}(y))}+F^{2}(b)\log\frac{F(b)}{F(v_{\tau}(y))}\right.\\ &\left.+[1-F(b)]^{2}\log\frac{1-F(v_{\tau}(y))}{1-F(b)}\right\},\end{aligned}$
@@
where $F(x)=1/[1+\exp(-x)]$ is the standard logistic CDF, $F_{a}^{b}$ is the standard logistic CDF truncated to $[a,b]$ with $-\infty<a<b<\infty$, and $F_{a,\mu,\sigma}^{b}(x)=F_{(a-\mu)/\sigma}^{(b-\mu)/\sigma}\left(\frac{x-\mu}{\sigma}\right)$ with location parameter $\mu\in\mathbb{R}$ and scale parameter $\sigma>0$.
@@ B.4 Some notes on the integral function $I(\alpha,\beta)$
In the derivation of the twCRPS for the log-normal distribution the integral function $I(\alpha,\beta)$ appears (equation(38)). It has the statistical interpretation as the probability $P(X>\alpha,Y<X+\beta)$ where $X$ and $Y$ are two independent standard normal random variables. The integral function $I$ can be expressed in terms of Owen’s $T$ function [Owen, 1956, 1980],
@@
Alternatively, we propose a fast and stable method for evaluating $I(\alpha,\beta)$ based on a recursive power series expansion, making use of (probabilist’s) Hermite polynomials ${\{H_{n}\}}_{n=0}^{\infty}$. The integral function $I$ has the Taylor series with respect to the second argument
@@
$C_{n}=$ $-\frac{\alpha\beta}{n}C_{n-1}-\frac{n-3}{(n-1)n}\beta^{2}C_{n-2},\qquad\qquad n\geq 3,$ (55)
$I_{n}=$ $C_{n}-\frac{n-2}{2(n-1)n}\beta^{2}I_{n-2},\qquad\qquad n\geq 3.$ (56)
@@
Appendix C Additional figures
Table 1: 80th and 90th percentile thresholds for each of the four clusters. The colors refer to Figure 1.
Cluster Cluster interpretation Number of locations 80th percentile 90th percentile
1 (purple) Inland locations 30 4.1m/s 5.7m/s
2 (green) More exposed inland locations 44 5.7m/s 7.2m/s
3 (red) Coastal locations 35 7.7m/s 9.8m/s
4 (blue) More exposed coastal locations 15 10.3m/s 12.3m/s
Figure 13: twCRPS skill for the CRPS-trained and twCRPS-trained EMOS model (truncated normal predictive distribution, 90th percentile threshold), compared to the raw ensemble for a range of evaluation thresholds between 0th and 99th percentile of the training distribution. Each line corresponds to one cluster. Note that due to the large percentage differences the relative improvement of the twCRPS model compared to the CRPS model is not very visible. The raw ensemble does not include the forecast control member as this member is not exchangeable with the rest of the ensemble.
@@ Appendix C Additional figures
0.8 0.7 -0.62 -0.43
0.8 -0.26 -0.36
0.85 0.14 -0.05
0.9 0.52 0.28
0.95 0.52 0.31
0.9 0.7 -3.34 -2.71
0.8 -1.93 -1.88
0.85 -0.97 -1.11
0.9 0.10 -0.22
0.95 0.27 -0.03
Figure 14: twCRPS skill score (over the CRPS-trained model) as a function of the evaluation threshold for twCRPS-trained models trained using different thresholds.
Figure 15: Brier skill score (over the CRPS-trained model) by cluster as a function of the evaluation threshold for twCRPS-trained models trained using different thresholds. Thresholds on the x-axis correspond to wind speeds between the 50th and 99th percentile of the training distribution. The vertical lines indicate the thresholds for which models were trained. Note the different x-axis starting points to Figure 6.